OpenStudy (abmon98):

As a car passes the point A on s straight road, its speed is 10 m/s. The car moves with constant acceleration a m/s² along the road for T seconds until it reaches the point B, where its speed is V m/s. The car travels at this speed for a further 10 seconds, when it reaches point C. From C it travels for a further T seconds with constant acceleration 3a m/a² until it reaches a speed of 20 m/s at the point D. Show that V= 125. Given that the distance between A and D is 675 m, Find the values of a and T.

ganeshie8 (ganeshie8):

maybe start by drawing Velocity-time graph

OpenStudy (abmon98):

x₁ = 0 + 10T + ½aT² v² = v₀²+2as vB² = 10²+ 2a(10T + ½aT²) vB = √[10²+ 2a(10T + ½aT²)] vC = VB = √[10²+ 2a(10T + ½aT²)] ◄─── Eq # 1 x₂ = x(B→C) = 10 * vB x₂ = 10 * √[10²+ 2a(10T + ½aT²)] v = v₀ + at 20 = vC + (3a)T vC = 20 - 3aT ◄─── Eq # 2

OpenStudy (abmon98):

solve to find T

OpenStudy (abmon98):

iam not that good on drawing graphs

ganeshie8 (ganeshie8):

okay, its not required

ganeshie8 (ganeshie8):

first lets try solving for T, using the fact that, time between A->B and C->D is same

OpenStudy (abmon98):

okay

ganeshie8 (ganeshie8):

scratch that, first lets try solving for \(\color{red}{V}\), using the fact that, time between A->B and C->D is same

ganeshie8 (ganeshie8):

A->B : V = 10 + aT ---------------(1) C->D : 20 = V + (3a)T --------------(2)

ganeshie8 (ganeshie8):

solve them both for V

OpenStudy (abmon98):

V=10 +aT 20=10+at+3at 10=4at t=10/4a t=5/2a

ganeshie8 (ganeshie8):

substitute that value of t in first equation, that gives u value of V

OpenStudy (abmon98):

V=10+a(5/2a) V=10+5a/2a V= 10+5/2 V=25

OpenStudy (abmon98):

sorry 12.5

ganeshie8 (ganeshie8):

yes

OpenStudy (abmon98):

Now we could find the distances am i right

ganeshie8 (ganeshie8):

oh wait, we need to find \(a\) and \(T\) still

OpenStudy (abmon98):

oh i forgot

ganeshie8 (ganeshie8):

u familiar wid below equation : v^2 - u^2 = 2*a*s ?

ganeshie8 (ganeshie8):

say, the distance between A->B is \(S1\), the distance between C->D is \(S2\) then, since the total distance is 625, \(S1 + 10*12.5 + S2 = 625\) ----------(3)

ganeshie8 (ganeshie8):

A->B : \(12.5^2 - 10^2 = 2*a*S1\) -----------(4) C->D : \(12.5^2 - 10^2 = 2*(2a)*S2\) --------(5)

ganeshie8 (ganeshie8):

eliminate \(a\) from equations (4) and (5)

ganeshie8 (ganeshie8):

then, u solve S1 and S2 distances

ganeshie8 (ganeshie8):

had we graphed Velocity-time graph, this wud have been much simpler to calculate :o

OpenStudy (abmon98):

what is a equal to

ganeshie8 (ganeshie8):

but ok, (3), (4), (5) are not looking that scary, u can solve them easily :)

ganeshie8 (ganeshie8):

eliminate \(a\) by dividing : (4) / (5)

OpenStudy (abmon98):

when i looked back at the book a=1/8

OpenStudy (abmon98):

how?

ganeshie8 (ganeshie8):

we wil wry about \(a\) later, first solve the distances \(S1\) and \(S2\)

ganeshie8 (ganeshie8):

(5)/ (4) : \(\large \frac{20^2-12.5^2}{12.5^2-10^2} = \frac{3S2}{S1}\) \(13S1 = 9S2\)------------(6)

ganeshie8 (ganeshie8):

u can solve (3) and (6) for \(S1\) and \(S2\) once u have distances, u can find \(a\) and \(T\)